# Muffled Murmurs: Environmental effects in the LISA stochastic signal from stellar-mass black hole binaries

### v1

@@ I Introduction
The detection of compact binary coalescences by the LIGO-Virgo-KAGRA (LVK) collaboration offers valuable insight into the sBBHs evolution, formation channels, and population properties [1, 2, 3, 4, 5]. GW190521 is an exceptional gravitational-wave (GW) event, with the primary black-hole (BH) mass residing in the mass gap predicted by pair-instability supernova theory, therefore challenging current astrophysical formation scenarios [6, 7]. One possible explanation is that this event occurred within the gaseous disk of an active galactic nucleus (AGN), where mass segregation and dynamical friction drive the migration of black holes near the disk’s center, enhancing merger rates and facilitating mass growth through repeated mergers and sustained accretion [8, 9, 10]. Therefore, probing environmental effects through gravitational-wave observations is essential for advancing our understanding of astrophysical processes.
@@
The Laser Interferometer Space Antenna (LISA) [11], a planned space-based gravitational-wave observatory operating in the mHz frequency band, is well suited for detecting the dynamical signatures induced by astrophysical environments [12, 13, 14, 15]. A primary reason for this is that environmental effects are typically more significant earlier in the inspiral. In the LISA band, extreme mass ratio inspirals and intermediate mass black hole binaries are typical probes of the strong dynamics induced by the astrophysical environment [16, 17, 18, 19, 20, 21, 22, 23, 24]. Yet, the resolvable sBBHs that form in gas-rich environments, e.g. in the disks of AGNs, are also potentially sensitive to environmental effects [25, 21, 26, 27]. However, the largest majority of sBBHs will not be detectable in the LISA band, resulting in the build-up of a SGWB [28, 29, 30].
Environmental effects typically induce additional energy dissipation that are pre-Newtonian or negative post-Newtonian (PN) relative to the leading point-particle GW flux [19]. Consequently, when the additional energy loss due to the environment dominates over the GW flux at low frequencies, there will be a significant drop in the SGWB relative to the vacuum case. A similar scenario has been proposed to explain pulsar timing array (PTA) measurements [31, 32, 33, 34, 35, 36], where the observed SGWB may originate from a population of supermassive binary black holes influenced by environmental effects [37, 38, 39].
@@
The main environmental effects expected to affect sBBHs in gas-rich environments are dynamical friction and accretion [19, 21, 25, 26]. Dynamical friction is the result of the gravitational interaction of each black hole with the density wake produced by its motion through a fluid, collisionless [40] or collisional [41, 42, 19]. Accretion affects the binary because the infalling gas carries energy and momentum, which are transferred to the black hole and change its mass and momentum [16, 21, 25]. Thus, we focus on the imprints of dynamical friction and accretion on the SGWB of sBBHs.
@@
Treating dynamical friction and accretion effects as perturbations on the Keplerian orbital motion, we derive their effect on the spectrum of the stochastic background from a population of sBBHs consistent with the observational constraints from LVK’s third observing run (O3) [43, 5]. We construct phenomenological parametric (and analytic) models for the SGWB with environmental effects, which can be readily used with the Bayesian tools developed in [44, 45] to quantify the detectability of accretion and dynamical friction with LISA observations.
@@
We find that the effect of gas accretion is not detectable in the LISA stochastic background, if the accretion rate is Eddington-limited, with the Eddington ration $f_{\rm Edd}\lesssim 10$. Meanwhile, dynamical friction from gas densities comparable to those expected in AGN disks would yield a measurable effect on the background’s spectrum. In more detail, LISA’s observations of the stochastic background will probe gas densities $\rho\gtrsim\mbox{a few}\,\times 10^{-10}\mathrm{g\,cm^{-3}}$, with densities larger than $10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$ detectable with large Bayes factors, even when accounting for the effect of the Galactic white-dwarf foreground. We also find that a sub-population of sBBHs undergoing dynamical friction in typical AGN disk gas densities can be probed, provided it contributes at least $\sim 10\%$ of the total SGWB, with the remainder arising from sBBHs in vacuum. Thus, this offers potential insight into the formation channels of sBBHs.
@@ II Stochastic gravitational-wave background from stellar binaries in vacuum
A Gaussian, isotropic, unpolarized, and stationary gravitational wave background (SGWB) is fully characterized by its spectral energy density, $\Omega_{\mathrm{GW}}(f)$, given by [53, 54, 43]
@@
[…]
where $\frac{dE_{\rm GW}}{df_{s}}(\bm{\phi})|_{f_{s}}$ is the source-frame energy spectrum radiated by a single source, evaluated at the source GW frequency $f_{s}=f(1+z)$ with $f$ being the detector frame GW frequency. In Eq. 2, the integration is performed over the distribution $p(\bm{\phi})$ of source parameters $\bm{\phi}$ (e.g. masses, spins, etc.). The quantity $R_{\mathrm{GW}}(z)$ is the comoving merger rate density of GW sources measured in the source frame, and $\mathcal{E}(z)=\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}$. We adopt the result of Planck18 [55] for the value of cosmology parameters.
For a binary with masses $m_{1},m_{2}$, to leading order in the PN expansion, the energy spectrum carried by gravitational waves emitted by a binary at a frequency $f_{s}$ is given by [54, 56]
@@
From Eqs. 2 and 3, the frequency-independent contribution to the SGWB can be absorbed into an overall amplitude $A_{\rm vac}$, yielding
@@
We adopt a standard astrophysical prescription for the merger rate and source population [5, 57, 58, 48, 59, 60]. The redshift-dependent merger rate is assumed to follow the cosmic star formation rate [57], weighted by metallicity [58] and convolved with a distribution of time delays [48]. The mass distribution follows the Power Law + Peak model [59] with negligible spins, consistent with LIGO-Virgo observations [61, 62]. With these models and the posterior distributions of their parameters inferred from Refs.[43, 5], we numerically evaluate Eq. 2 via Monte Carlo integration to generate posterior predictions for the SGWB. The median SGWB posterior prediction and the corresponding model parameters are adopted as fiducial values. A comparison with the O3-based forecast is shown in Appendix A, demonstrating consistency across the relevant frequency range.
@@ III.1 Dynamical friction
where $f_{A}^{*}=50v_{s}/(11\pi m_{A})$. From the energy-balance law, the rate of change of binding energy $E_{\rm b}$ is given by $\dot{E}_{\rm b}=-\dot{E}_{\rm GW}-\dot{E}_{\rm DF}$, yielding an additional contribution to $\dot{f}$:
@@
×$\times\left.\left[m_{1}^{3}\ln\left(\dfrac{f_{1}^{*}}{f_{s}}\right)+(m_{1}\leftrightarrow m_{2})\right]\right\}^{-1},$ (10)
where $(dE_{\rm GW}/df_{s})_{\rm vac}$ is given by Eq. 3. The energy flux due to dynamical friction is a -5.5PN relative to the GW flux, owing to the $f_{s}^{-11/3}$ scaling. Consequently, at low frequencies dynamical friction will dominate the orbital evolution and energy loss, and can deplete the SGWB.
At a critical turning point frequency $f_{\rm turn,DF}$, the flux due to dynamical friction becomes comparable to that of GW emission. For $f_{s}<f_{\rm turn,DF}$, the amplitude of the GW spectrum will start to decrease as more energy flows through the dynamical friction channel than the GW emission channel. We compute $f_{\rm turn,DF}$ from $|\dot{E}_{\rm DF}|=|\dot{E}_{\rm GW}|$, which amounts to setting the $\rho$-dependent term in the denominator of Eq. 10 to 1.
To gain insight onto $f_{\rm turn,DF}$, let us note that typical astrophysical systems yield $f_{A,*}\sim 10^{-1}$z, much higher than the observed frequencies that we want to probe. Observing that $f_{A,*}\gg f_{\rm turn,DF}$ and focusing on nearly equal mass systems, we obtain
@@
×$\times\left(\dfrac{\rho}{10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}}\right)^{3/11},$ (11)
@@
In Fig. 1, we show the SGWB computed with Eq. 10. As for the vacuum case, we perform a three dimensional Monte Carlo population integral. SGWB are bracketed by a few reference $\rho$ values. As $\rho$ is increased, the turning point in the spectrum does increase as expected from Eq. 11. Further, the spectral index of the SGWB asymptotes to a value of $13/3$ at low frequencies and $2/3$ at high frequencies, the latter being the vacuum result. For typical disk densities of $\rho\sim 10^{-11}\mathrm{g}\,\mathrm{cm}^{-3}$$\rho\sim 10^{-8}\mathrm{g}\,\mathrm{cm}^{-3}$, the stochastic signal typically lies above the power-law sensitivity (PLS) and Bayesian power-law sensitivity (BPLS) curves [45] for $f\gtrsim 10^{-3}$z. We note that both the PLS and BPLS curves assume power-law SGWB signals, with the signal-to-noise ratio (SNR) and Bayes factor quantifying signal detectability relative to noise, and are included here merely as references. Additionally, for the typical disk densities considered here, we can expect that the dynamical friction effects are potentially measurable since the turning point occurs in the sensitive part of the LISA band. Thus, the SGWB of sBBHs is potentially detectable and the effects of dynamical friction are also potentially measurable. However, to rigorously determine the detectability of the signal and measurability of the parameters, we perform a detailed Bayesian analysis in Sec. IV.
@@ III.2 Gas accretion
The gas from the surrounding disk will accrete onto the two black holes. We model the accretion of the $A$-th body as $\dot{m}_{A,\rm Edd}=L_{A,\rm Edd}/\zeta$, where $L_{A,\rm Edd}$ is the Eddington luminosity and $\zeta$ is the radiative efficiency [19]. We pick a conservative value for the radiative efficiency of $\zeta=0.1$ and the resulting accretion rate is $\dot{m}_{A,\rm Edd}\simeq 2.2\times 10^{-8}(m_{A}/M_{\odot})M_{\odot}\mathrm{yr}^{-1}$ [19, 21]. For simplicity, we parameterize the accretion rate of both black holes by the same Eddington ratio $f_{\rm Edd}\equiv\dot{m}_{A}/\dot{m}_{A,\rm Edd}$. Doing so, the mass as a function of time reads [21]
@@
One could also obtain the same back-reaction using the energy-balance law but after accounting for the time-variation of the Hamiltonian under the adiabatic mass increase. We explicitly show the equivalence of the two balance laws in Appendix B.2, as an incorrect version of the energy balance law has been used in the literature [67, 68, 52, 49] to obtain the back-reaction for a time-dependent Hamiltonian. The total back-reaction due to accretion and GW emission is then the sum of Eqs. 14 and 4. We also emphasize that Eq. 4 is evaluated with the initial masses, i.e. we neglect contributions due to cross terms between accretion and radiation reaction because we effectively treat these effects to leading order in a multiple-timescale analysis [69].
In this work, we do not consider relativistic corrections to the hydrodynamical drag (such as in [16, 18]) as (i) they are more important for EMRIs than for comparable mass stellar binaries, and (ii) we are primarily interested in how accretion effects contribute as an additional energy dissipation channel, for which the non-relativistic treatment is sufficient. We comment further on the validity of the accretion modeling in Appendix B. Thus, using Eq. 14 together with Eqs. 4 and 3, we obtain the energy spectrum of gravitational waves in the presence of accretion as
@@
×$\times\left[1+\dfrac{5(f_{\mathrm{Edd}}/\tau)(5+3\xi)}{96\pi^{8/3}m_{0}^{5/3}\eta}f_{s}^{-8/3}\right]^{-1},$ (15)
where $(dE_{\rm GW}/df_{s})_{\rm vac}$ is given by Eq. 3 and evaluated with the initial value of the masses.
Figure 2: The SGWB spectra $\Omega_{\mathrm{GW}}(f)$, driven by gas accretion, for a range of Eddington ratios $f_{\mathrm{Edd}}$. The solid lines correspond to different values of $f_{\rm Edd}$, with the vacuum case corresponding to $f_{\rm Edd}=0$. The gray dashed and black dotted curves indicate the LISA power-law sensitivity (PLS) and Bayesian power-law sensitivity (BPLS) under the same assumptions made in Sec. III.1.
@@
×$\quad\times\left(\frac{5+3\xi}{8}\right)^{3/8}\left(\frac{\tau}{4.5\times 10^{7}\,\mathrm{yr}}\right)^{-3/8}\,.$ (16)
In addition, as the total mass increases, the turning point shifts toward lower frequencies. For typical values of $f_{\rm Edd}\sim 1$, we find that $f_{\rm turn}\sim 10^{-4}$z, implying that the effects are significant only at the lower end of the LISA sensitivity band. Only with much higher values of $f_{\rm Edd}\sim 10^{3}$, which may not be astrophysically realistic, we have $f_{\rm turn}\sim 10^{-3}$z, which is comparable to that of dynamical friction. As expected, the effect of gas accretion occurs at -4PN order. The key reason that gas accretion has an overall lower turning point than dynamical friction is its higher (less negative) PN order relative to GW emission. [^1] Typically when a particular effect scales as $v^{-n}$ relative to GW emission, with $v\equiv(\pi mf)^{1/3}$ and $n>0$, the effect is stronger for larger $n$ due to $v\ll 1$..
In Fig. 2, we show the SGWB computed with Eq. 15 for different astrophysical values of $f_{\rm Edd}$, and we set $\xi=1$ for simplicity. As expected, the signal is suppressed at low frequencies due to the additional energy loss channel, and asymptotes to the vacuum SGWB at high frequencies. Although the stochastic signals shown in Fig. 2 lie above the PLS and BPLS curves for the frequency range of $f\gtrsim 10^{-3}$z, suggesting that they are detectable, the turning points are outside the LISA SGWB sensitivity band for the range of astrophysical $f_{\rm Edd}$ values considered here. Thus, we do not expect to clearly measure $f_{\rm Edd}$ with LISA and distinguish gas accretion from vacuum. As with the case of dynamical friction, this requires a more rigorous Bayesian analysis, which we do in Sec. IV.

### v2

@@ I Introduction
The detection of compact binary coalescences by the LIGO-Virgo-KAGRA (LVK) collaboration offers valuable insight into the stellar-mass black hole binaries (sBBHs) evolution, formation channels, and population properties [1, 2, 3, 4, 5]. GW190521 is an exceptional gravitational-wave (GW) event, with the primary black-hole (BH) mass residing in the mass gap predicted by pair-instability supernova theory, therefore challenging current astrophysical formation scenarios [6, 7]. One possible explanation is that this event occurred within the gaseous disk of an active galactic nucleus (AGN), where mass segregation and dynamical friction drive the migration of black holes near the disk’s center, enhancing merger rates and facilitating mass growth through repeated mergers and sustained accretion [8, 9, 10]. Therefore, probing environmental effects through gravitational-wave observations is essential for advancing our understanding of astrophysical processes.
@@
The Laser Interferometer Space Antenna (LISA) [11], a planned space-based gravitational-wave observatory operating in the mHz frequency band, is well suited for detecting the dynamical signatures induced by astrophysical environments [12, 13, 14, 15]. A primary reason for this is that environmental effects are typically more significant earlier in the inspiral. In the LISA band, extreme mass ratio inspirals and intermediate mass black hole binaries are typical probes of the strong dynamics induced by the astrophysical environment [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Yet, the resolvable sBBHs that form in gas-rich environments, e.g. in the disks of AGNs, are also potentially sensitive to environmental effects [26, 22, 27, 28]. However, the largest majority of sBBHs will not be detectable in the LISA band, resulting in the build-up of a stochastic gravitational-wave background (SGWB) [29, 30, 31]. In general, a SGWB is a superposition of GW signals from an unresolved population of sources [32, 33, 34]. Computing the SGWB from unresolved sBBHs in the LISA band has been recently explored with the assumption that the the binaries are in vacuum [31, 35, 36, 37]. Since the amplitude of the vacuum SGWB from sBBHs may be measurable at the percent level [31], we extend previous vacuum-based studies by investigating how environmental effects could imprint observable deviations.
@@
Environmental effects typically induce additional energy dissipation that can dominate over the GW flux at low frequencies [19, 20], thereby suppressing the SGWB relative to the vacuum case. A similar scenario has been proposed to explain pulsar timing array (PTA) measurements [38, 39, 40, 41, 42, 43], where the observed SGWB may originate from a population of supermassive binary black holes influenced by environmental effects [44, 45, 46, 47]. The main environmental effects expected to affect sBBHs in gas-rich environments are dynamical friction and accretion [20, 22, 26, 27] (see also [19] for an overview of other environmental effects). Dynamical friction is the result of the gravitational interaction of each black hole with the density wake produced by its motion through a fluid, collisionless [48] or collisional [49, 50, 20]. Accretion affects the binary because the infalling gas carries energy and momentum, which are transferred to the black hole and change its mass and momentum [16, 22, 26]. Thus, we focus on the imprints of dynamical friction and accretion on the SGWB of sBBHs.
@@
Treating dynamical friction and accretion effects as perturbations on the Keplerian orbital motion, we derive their effect on the spectrum of the stochastic background from a population of sBBHs consistent with the observational constraints from LVK’s third observing run (O3) [51, 5]. We construct phenomenological parametric models for the SGWB with environmental effects, which can be readily used with the Bayesian tools developed in [52, 53] to quantify the detectability of accretion and dynamical friction with LISA observations.
@@
We find that the effect of gas accretion is not detectable in the LISA stochastic background, if the accretion rate is Eddington-limited, with the Eddington rate $f_{\rm Edd}\lesssim 10$. Meanwhile, dynamical friction from gas densities comparable to those expected in AGN disks would yield a measurable effect on the background’s spectrum. In more detail, LISA’s observations of the stochastic background will probe gas densities $\rho\gtrsim\mbox{a few}\,\times 10^{-10}\mathrm{g\,cm^{-3}}$, with densities larger than $10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$ detectable with large Bayes factors, even when accounting for the effect of the Galactic white-dwarf foreground. We also find that a sub-population of sBBHs undergoing dynamical friction in typical AGN disk gas densities can be probed. Thus, this may offer potential insight into the formation channels of sBBHs.
@@ II Stochastic gravitational-wave background from stellar binaries in vacuum
A Gaussian, isotropic, unpolarized, and SGWB is fully characterized by its spectral energy density, $\Omega_{\mathrm{GW}}(f)$, given by [32, 33, 51]
@@
$\Omega_{\mathrm{GW}}(f)=\frac{f}{\rho_{c}H_{0}}$ $\iint dzd\bm{\phi}\frac{R_{\mathrm{GW}}(z)}{(1+z)\mathcal{E}(z)}p(\bm{\phi})\left.\frac{dE_{\rm GW}}{df_{s}}(\bm{\phi})\right|_{f_{s}},$ (2)
where $\frac{dE_{\rm GW}}{df_{s}}(\bm{\phi})|_{f_{s}}$ is the source-frame energy spectrum radiated by a single source, evaluated at the source GW frequency $f_{s}=f(1+z)$ with $f$ being the detector frame GW frequency. In Eq.˜2, the integration is performed over the distribution $p(\bm{\phi})$ of source parameters $\bm{\phi}$ (e.g. masses, spins, etc.). The quantity $R_{\mathrm{GW}}(z)$ is the comoving merger rate density of GW sources measured in the source frame, and $\mathcal{E}(z)=\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}$. We adopt the result of Planck18 [61] for the value of cosmology parameters.
For a binary with masses $m_{1},m_{2}$, to leading order in the post-Newtonian (PN) expansion [^1] In PN theory one expands in the weak-field, slow-motion limit. An nnPN term scales as $v^{2n}$ relative to the leading contribution [34], with $v=(\pi mf_{s})^{1/3}$ being the orbital velocity. , the energy spectrum carried by gravitational waves emitted by a binary at a frequency $f_{s}$ is given by [33, 34]
@@
From Eqs.˜2 and 3, the frequency-independent contribution to the SGWB can be absorbed into an overall amplitude $A_{\rm vac}$, yielding
@@
We adopt a standard astrophysical prescription for the merger rate and source population [5, 62, 63, 56, 64, 65]. The redshift-dependent merger rate is assumed to follow the cosmic star formation rate [62], weighted by metallicity [63] and convolved with a distribution of time delays [56]. The mass distribution follows the Power Law + Peak model [64] with negligible spins, consistent with LIGO-Virgo observations [66, 67]. With these models and the posterior distributions of their parameters inferred from Refs.[51, 5], we numerically evaluate Eq.˜2 via Monte Carlo integration to generate posterior predictions for the SGWB. The median SGWB posterior prediction and the corresponding model parameters are adopted as fiducial values. A comparison with the O3-based forecast is shown in Appendix A, demonstrating consistency across the relevant frequency range. Further, we note that one can alternatively use the chirp rate $\dot{f_{s}}$ to characterize the binary evolution and compute the number of systems within a given GW frequency bin of interest [68]. In our approach, however, the inclusion of additional environmental dissipative channel modifies the spectral shape through its impact on the energy spectrum, while the overall normalization remains fixed by the merger rate informed by LVK population constraints, rather than by renormalizing the number of sources per frequency bin.
@@ III.1 Dynamical friction
where $f_{A}^{*}=50v_{s}/(11\pi m_{A})$ and the shorthand $(m_{1}\leftrightarrow m_{2})$ denotes the same term with $m_{1}$ and $m_{2}$ interchanged. From the energy-balance law, the rate of change of binding energy $E_{\rm b}$ is given by $\dot{E}_{\rm b}=-\dot{E}_{\rm GW}-\dot{E}_{\rm DF}$, yielding an additional contribution to $\dot{f}$:
@@
$\times\left.\left[m_{1}^{3}\ln\left(\dfrac{f_{1}^{*}}{f_{s}}\right)+(m_{1}\leftrightarrow m_{2})\right]\right\}^{-1},$ (10)
where $(dE_{\rm GW}/df_{s})_{\rm vac}$ is given by Eq.˜3. The energy flux due to dynamical friction is a -5.5PN relative to the GW flux, owing to the $f_{s}^{-11/3}$ scaling. Consequently, at low frequencies dynamical friction will dominate the orbital evolution and energy loss, and can deplete the SGWB.
At a critical turning point frequency $f_{\rm turn,DF}$, the flux due to dynamical friction becomes comparable to that of GW emission. For $f_{s}<f_{\rm turn,DF}$, the amplitude of the GW spectrum will start to decrease as more energy flows through the dynamical friction channel than the GW emission channel. We compute $f_{\rm turn,DF}$ from $|\dot{E}_{\rm DF}|=|\dot{E}_{\rm GW}|$, which amounts to setting the $\rho$-dependent term in the denominator of Eq.˜10 to 1.
To gain insight onto $f_{\rm turn,DF}$, let us note that typical astrophysical systems yield $f_{A,*}\sim 10^{-1}$Hz, much higher than the observed frequencies that we want to probe. Observing that $f_{A,*}\gg f_{\rm turn,DF}$ and focusing on nearly equal mass systems, we obtain
@@
$\times\left(\dfrac{\rho}{10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}}\right)^{3/11},$ (11)
@@
In Fig.˜1, we show the SGWB computed with Eq.˜10. As for the vacuum case, we perform a three dimensional Monte Carlo population integral. SGWB are bracketed by a few reference $\rho$ values. As $\rho$ is increased, the turning point in the spectrum does increase as expected from Eq.˜11. Further, the spectral index of the SGWB asymptotes to a value of $13/3$ at low frequencies and $2/3$ at high frequencies, the latter being the vacuum result. For typical disk densities of $\rho\sim 10^{-11}\mathrm{g}\,\mathrm{cm}^{-3}$$\rho\sim 10^{-8}\mathrm{g}\,\mathrm{cm}^{-3}$, the stochastic signal typically lies above the power-law sensitivity (PLS) and Bayesian power-law sensitivity (BPLS) curves [53] for $f\gtrsim 10^{-3}$Hz. We note that both the PLS and BPLS curves assume power-law SGWB signals, with the signal-to-noise ratio (SNR) and Bayes factor quantifying signal detectability relative to noise, and are included here merely as references. Additionally, for the typical disk densities considered here, we can expect that the dynamical friction effects are potentially measurable since the turning point occurs in the sensitive part of the LISA band. Thus, the SGWB of sBBHs is potentially detectable and the effects of dynamical friction are also potentially measurable. However, to rigorously determine the detectability of the signal and measurability of the parameters, we perform a detailed Bayesian analysis in Sec.˜IV.
@@ III.2 Gas accretion
The gas from the surrounding disk will accrete onto the two black holes. We model the accretion of the $A$-th body as $\dot{m}_{A,\rm Edd}=L_{A,\rm Edd}/\zeta$, where $L_{A,\rm Edd}$ is the Eddington luminosity and $\zeta$ is the radiative efficiency [20] (see also [72] for a detailed review of accretion modeling). We pick a conservative value for the radiative efficiency of $\zeta=0.1$ and the resulting accretion rate is $\dot{m}_{A,\rm Edd}\simeq 2.2\times 10^{-8}(m_{A}/M_{\odot})M_{\odot}\mathrm{yr}^{-1}$ [20, 22]. For simplicity, we parameterize the accretion rate of both black holes by the same Eddington ratio $f_{\rm Edd}\equiv\dot{m}_{A}/\dot{m}_{A,\rm Edd}$. Doing so, the mass as a function of time reads [22]
@@
One could also obtain the same back-reaction using the energy-balance law but after accounting for the time-variation of the Hamiltonian under the adiabatic mass increase. We explicitly show the equivalence of the two balance laws in Appendix B.2, as an incorrect version of the energy balance law has been used in the literature [74, 75, 60, 57] to obtain the back-reaction for a time-dependent Hamiltonian. The total back-reaction due to accretion and GW emission is then the sum of Eqs.˜14 and 4. We also emphasize that Eq.˜4 is evaluated with the initial masses, i.e. we neglect contributions due to cross terms between accretion and radiation reaction because we effectively treat these effects to leading order in a multiple-timescale analysis [76].
In this work, we do not consider relativistic corrections to the hydrodynamical drag (such as in [16, 18]) as (i) they are more important for extreme-mass-ratio inspirals than for comparable mass stellar binaries, and (ii) we are primarily interested in how accretion effects contribute as an additional energy dissipation channel, for which the non-relativistic treatment is sufficient. We comment further on the validity of the accretion modeling in Appendix B. Thus, using Eq.˜14 together with Eqs.˜4 and 3, we obtain the energy spectrum of gravitational waves in the presence of accretion as
@@
$\times\left[1+\dfrac{5(f_{\mathrm{Edd}}/\tau)(5+3\xi)}{96\pi^{8/3}m_{0}^{5/3}\eta}f_{s}^{-8/3}\right]^{-1},$ (15)
where $(dE_{\rm GW}/df_{s})_{\rm vac}$ is given by Eq.˜3 and evaluated with the initial value of the masses.
Figure 2: The SGWB spectra $\Omega_{\mathrm{GW}}(f)$, driven by gas accretion, for a range of Eddington ratios $f_{\mathrm{Edd}}$. The solid lines correspond to different values of $f_{\rm Edd}$, with the vacuum case corresponding to $f_{\rm Edd}=0$. The gray dashed and black dotted curves indicate the LISA power-law sensitivity (PLS) and Bayesian power-law sensitivity (BPLS) under the same assumptions made in Sec.˜III.1.
@@
$\quad\times\left(\frac{5+3\xi}{8}\right)^{3/8}\left(\frac{\tau}{4.5\times 10^{7}\,\mathrm{yr}}\right)^{-3/8}\,.$ (16)
In addition, as the total mass increases, the turning point shifts toward lower frequencies. For typical values of $f_{\rm Edd}\sim 1$, we find that $f_{\rm turn}\sim 10^{-4}$Hz, implying that the effects are significant only at the lower end of the LISA sensitivity band. Only with much higher values of $f_{\rm Edd}\sim 10^{3}$, which may not be astrophysically realistic, we have $f_{\rm turn}\sim 10^{-3}$Hz, which is comparable to that of dynamical friction. As expected, the effect of gas accretion occurs at -4PN order due to the $f_{s}^{-8/3}$ scaling. The key reason that gas accretion has an overall lower turning point than dynamical friction is its higher (less negative) PN order relative to GW emission. [^2] Typically when a particular effect scales as $v^{-n}$ relative to GW emission, with $v=(\pi mf_{s})^{1/3}$ and $n>0$, the effect is stronger for larger $n$ due to $v\ll 1$..
In Fig.˜2, we show the SGWB computed with Eq.˜15 for different astrophysical values of $f_{\rm Edd}$, and we set $\xi=1$ for simplicity. As expected, the signal is suppressed at low frequencies due to the additional energy loss channel, and asymptotes to the vacuum SGWB at high frequencies. Although the stochastic signals shown in Fig.˜2 lie above the PLS and BPLS curves for the frequency range of $f\gtrsim 10^{-3}$Hz, suggesting that they are detectable, the turning points are outside the LISA SGWB sensitivity band for the range of astrophysical $f_{\rm Edd}$ values considered here. Thus, we do not expect to clearly measure $f_{\rm Edd}$ with LISA and distinguish gas accretion from vacuum. As with the case of dynamical friction, this requires a more rigorous Bayesian analysis, which we do in Sec.˜IV.